# thanks: Haihao Shi and Zhenyang Huang contribute equally to this workthanks: Corresponding authorthanks: Corresponding author

### v1

@@ Abstract
Neutrinos can experience fast flavor conversions (FFCs) in highly dense astrophysical environments, such as core-collapse supernovae and neutron star mergers, potentially affecting energy transport and other processes. The simulation of fast flavor conversions under realistic astrophysical conditions requires substantial computational resources and involves significant analytical challenges. While machine learning methods like Multilayer Perceptrons have been used to accurately predict the asymptotic outcomes of FFCs, their ‘black-box’ nature limits the extraction of direct physical insight. To address this limitation, we employ two distinct interpretable machine learning frameworks—Kolmogorov-Arnold Networks (KANs) and Sparse Identification of Nonlinear Dynamics (SINDy)—to derive physically insightful models from a FFC simulation dataset. Our analysis reveals a fundamental trade-off between predictive accuracy and model simplicity. The KANs demonstrates high fidelity in reconstructing post-conversion neutrino energy spectra, achieving accuracies of up to 90%. In contrast, SINDy uncovers a remarkably concise, low-rank set of governing equations, offering maximum interpretability but with lower predictive accuracy. Critically, by analyzing the interpretable model, we identify the number density of heavy-lepton neutrinos as the most dominant factor in the system’s evolution. Ultimately, this work provides a methodological framework for interpretable machine learning that supports genuine data-driven physical discovery in astronomy and astrophysics, going beyond prediction alone.
@@ I Introduction
A significant recent advancement in this area is the discovery of fast flavor conversions (FFCs), which occur over remarkably short spatial scales in the vacuum [18, 19, 20, 21, 22, 23]. A necessary and sufficient condition for the occurrence of FFC is that the angular distribution of the neutrino lepton number, defined as:
$G(\mathbf{v})=\frac{\sqrt{2}G_{\text{F}}}{(2\pi)^{3}}\int E_{\nu}^{2}dE_{\nu}\Big{[}(f_{\nu_{e}}(p)-f_{\nu_{x}}(p))$ (1)
$-(f_{\bar{\nu}_{e}}(p)-f_{\bar{\nu}_{x}}(p))\Big{]},$
crosses zero at some $\mathbf{v}=\mathbf{v}(\mu,\phi_{\nu})$, where $\mu=\cos\theta_{\nu}$ [24]. Here, $G_{\rm{F}}$ denotes the Fermi coupling constant, while $E_{\nu}$, $\theta_{\nu}$, and $\phi_{\nu}$ represent the neutrino energy, and the zenith and azimuthal angles of the neutrino velocity, respectively. The $f_{\nu}$’s denote neutrino occupation numbers for different flavors, with $\nu_{x}$ and $\bar{\nu}_{x}$ referring to heavy-lepton neutrinos and antineutrinos. Following state-of-the-art CCSNe and NSM simulations, we assume $\nu_{x}$ and $\bar{\nu}_{x}$ share similar angular distributions. Under this assumption, Equation 2 simplifies to the conventional definition of the neutrino electron lepton number ($\nu$ELN).
The extremely short characteristic scales of FFC, in stark contrast to the much coarser resolution of typical hydrodynamic simulations of CCSNe and NSMs, present a substantial challenge for seamless integration. The vast disparity in temporal and spatial scales across the various physical processes further renders direct, first-principles simulations computationally prohibitive. Compounding these difficulties, certain simulation details also require careful attention—for instance, while the equipartition condition appears reasonable for out-going neutrinos, significant deviations arise for incoming neutrinos, affecting the matter profiles [25]. A promising strategy is to conduct small-scale FFCs simulations and apply the insights gained to inform larger hydrodynamic models [26, 27, 28, 29, 30, 31, 32]. To address this, extensive research has focused on local dynamical simulations with periodic boundary conditions [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], where quasistationary flavor states have been observed in the neutrino gas and are well described by analytical models [44]. Recent studies suggest that when the initial symmetry of the neutrino field is broken, the nature of flavor conversion changes, potentially offering deeper insights into FFCs [45, 23, 46, 47]. On the theoretical side, a quantum framework describing neutrino interactions with flavor wave quanta has recently been proposed [48]. On the computational side, an approximate method to predict the asymptotic state of FFCs without solving the full quantum kinetic equations in the three-flavor framework has also been developed [49]. These advances offer potential solutions to the computational challenges of fast flavor evolution.
@@
Although flavor equilibration may occur immediately following rapid FFCs, this is not generally the asymptotic outcome, as feedback from subsequent collisional neutrino-matter interactions becomes increasingly important—particularly when FFCs occur inside the neutrinosphere [50]. Effective analytical methods have been developed to enable efficient and robust modeling of collisional feedback within the neutrinosphere [51]. Nonetheless, leveraging artificial intelligence to uncover hidden physical patterns in data remains a highly promising direction. This is because traditional analytical approaches typically involve specific assumptions and may not fully capture all aspects of the system’s complexity, creating unique opportunities for data-driven methods. With the rapid development of various neutrino telescopes, employing AI techniques offers an effective way to handle the vast volumes of observational data expected in the future and to extract the underlying physical insights. Recently, advances in machine learning, particularly neural networks, have been increasingly applied to astrophysics, including cosmological simulations [52, 53, 54, 55, 56] and astronomical data processing [57, 58, 59, 60, 61, 62, 63, 64, 65]. Recent studies have shown that artificial neural networks (NNs) can accurately predict the asymptotic outcomes of FFCs in both single- and multi-energy neutrino gases within the moments framework [66, 67, 68].
While deep learning has achieved remarkable success in certain research areas, many networks are parameter-intensive and lack transparency, making their internal mechanisms difficult to interpret. This remains a key challenge in AI-driven scientific research [69]. Addressing this issue requires moving beyond traditional “black box” models, as a deeper understanding of phenomena demands interpretability [70]. Interpretable networks are crucial for processing and analyzing data that may contain FFCs signatures. They not only enhance efficiency but also facilitate the extraction of physical features by tracing the model’s decision rationale and other similar methods, potentially leading to new hypotheses and insights in astronomy and physics.
Recently, Kolmogorov-Arnold Networks (KANs) have emerged as a promising solution. Inspired by the Kolmogorov-Arnold representation theorem, KANs introduce learnable activation functions that significantly improve model expressiveness while maintaining interpretability [71]. These advances open new possibilities for reanalyzing astronomical data and uncovering hidden physical features through interpretable modeling. Compared to conventional neural networks that focus on extracting features from data, KANs are more inclined to discover underlying physical laws. In parallel, the Sparse Identification of Nonlinear Dynamics (SINDy) framework has demonstrated effectiveness in recovering compact, symbolic representations of governing equations from data [72]. By promoting sparsity in the candidate function space, SINDy identifies the minimal set of terms required to describe system dynamics, offering a powerful tool for discovering interpretable physical laws from simulation outputs or observations.
In section II, we briefly introduce the structure of the model input. section III details the construction of KANs. In section IV, we describe the working principles of SINDy and how it is applied to the FFCs dataset. The reconstruction performance of KANs and SINDy for post-FFC supernova neutrino spectra is demonstrated in section V. Subsequently, section VI presents the extraction of physical insights from both the trained KANs and SINDy models. In section VII, we discuss the trade-off between predictive accuracy and interpretability in KAN and PySINDy. Finally, section VIII provides a summary of this work.
II Model Input
We first briefly introduce the input dataset used in this work, which is the same as that in [67]. This FFCs simulates a multienergy and multiangle neutrino gas in a 1D box, following the framework in [36]. For training and testing our models, we construct a dataset combining maximum entropy and Gaussian initial neutrino angular distributions. The final outcome of FFCs is determined by the three-flavor survival probability:
$P_{\text{sur}}(\mu)=\begin{cases}\frac{1}{3},&\text{for }\mu<\mu_{c}\\ \mathcal{S}_{3}(\mu),&\text{for }\mu>\mu_{c}\end{cases},$ (3)
where $\mathcal{S}_{3}(\mu)=1-\frac{2}{3}h\left(\left|\mu-\mu_{c}\right|/\zeta\right)$,$\quad h(x)=\left(x^{2}+1\right)^{-1/2}$. The parameter $\zeta$ is chosen to ensure the continuity of $P_{\text{sur}}(\mu)$, and $\mu^{<}$, $\mu^{>}$ are defined based on the comparison of the integrals:
@@ II Model Input
where $\Theta$ denotes the Heaviside theta function, $G(\mu)=\int_{0}^{2\pi}\mathrm{~{}d}\phi_{\nu}G(\mathbf{v})$ [44, 66].
The model parameters are as follows: $\alpha\in(0,2.5)$, $F_{\nu_{x},(i)}\in(0,1)$, $\bar{F}_{\nu_{e},(i)}\in(0.4F_{\nu_{x},(i)},F_{\nu_{x},(i)})$, and $F_{\nu_{e},(i)}\in(0.4\bar{F}_{\nu_{e},(i)},\bar{F}_{\nu_{e},(i)})$, which follows the expected hierarchy: $F_{\nu_{e}}\lesssim\bar{F}_{\nu_{e}}\lesssim F_{\nu_{x}}$. The neutrino number densities $n_{\nu,i}$ are sampled from a half-normal distribution with a mean of zero and a standard deviation of $0.1n_{\nu}$, where $n_{\nu}$ is the energy-integrated neutrino number density. With $F_{\nu}=(I_{1}/I_{0})_{\nu}$, the quantities with and without the subscript $i$ refer to the energy-integrated spectrum and specific energy bin, respectively. The neutrino number densities in a given energy bin must be smaller than the corresponding energy-integrated values. This is consistent with [67].
@@
Our 11-dimensional input vector consists of two sets of features. The first nine quantities primarily describe the moments of the neutrino distribution, as defined in Table 1. In addition, we include two energy-dependent components, $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$, defined for each bin $i$ as follows [67]:
[…]
with an identical expression for antineutrinos. In practice, each $E_{\nu,i}$ is computed from the ratio $F_{\nu,i}/F_{\nu}^{\rm tot}$ and then clipped to the interval [0, 60]060[0,\,60][ 0 , 60 ] to ensure numerical stability. A mask is applied to retain only those bins for which both $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$ are nonnegative; this masking is applied consistently to both inputs and targets so that the original training‐set composition is preserved. Finally, $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$ are concatenated (in that order) to the nine moment‐integrated inputs, yielding an 11‐dimensional vector for each sample. Throughout this process, the input dataset remains unchanged except for the bin‐mask filtering, thereby ensuring that the models learn from a physically consistent input distribution.
@@ III Kolmogorov-Arnold Networks
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$ (6)
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{KANs}(x)=(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{0})(x).$ (8)
@@
Figure 1 shows the training metrics through the training process, in which the calculation methods for the two error metrics are presented. Both metrics are forms of the Mean Absolute Error (MAE), calculated on the validation set. A small constant, $\epsilon=10^{-8}$, is added to the denominator in both cases to ensure numerical stability.
The overall relative error (represented by the blue line in the figure), denoted as $\mathcal{L}_{\text{rel}}$, measures the average relative error across all eight output neurons of the model. For a validation set with $N$ samples, where $\hat{y}_{j,i}$ is the model’s $i$-th predicted output for the $j$-th sample and $y_{j,i}$ is the corresponding true value, it is calculated as:
The $\bm{\nu_{e}+\bar{\nu}_{e}}$ error (represented by the red line), denoted as $\mathcal{L}_{\nu}$, is designed to assess the model’s performance on a key parameter. It specifically evaluates the model’s ability to predict the total contribution of electron neutrinos and antineutrinos. For each sample $j$, we define $S_{\text{pred},j}$ and $S_{\text{true},j}$ as the predicted and true sums for this total neutrino contribution, respectively. The error $\mathcal{L}_{\nu}$ is then the MAE calculated between these summed scalar values, averaged across all samples in the validation set:
Figure 2: Visualization of the trained KAN architecture used to model the fast flavor conversion process. The eleven input neurons at the bottom correspond, from left to right, to the physical features of the pre-conversion neutrino gas as detailed in Equation 4. The eight output neurons at the top yield the predicted post-conversion radial moments for all relevant flavors ($I_{0}^{\nu_{e}},I_{1}^{\nu_{e}},I_{0}^{\bar{\nu}_{e}},I_{1}^{\bar{\nu}_{e}},I_{0}^{\nu_{x}},I_{1}^{\nu_{x}},I_{0}^{\bar{\nu}_{x}},I_{1}^{\bar{\nu}_{x}}$, from left to right). The curves on the edges represent the learned one-dimensional activation functions, which are parameterized by B-splines. The activation shapes shown are taken from a representative training run, illustrating the complex functional relationships captured by the model.
@@
However, we identified that the symbolic regression process, a key step for KAN interpretability, exhibits a notable sensitivity to the random seed used in training. To obtain a robust and canonical symbolic model, we adopted an ensemble averaging approach. Specifically, the network was trained multiple times with different random seeds, and the B-spline coefficients for each learnable activation function were averaged across all runs. By performing symbolic regression on this stable set of averaged parameters, we mitigate the effects of stochastic initialization and derive a more representative model of the underlying dynamics.

### v2

@@ Abstract
Neutrinos can experience fast flavor conversions (FFCs) in highly dense astrophysical environments, such as core-collapse supernovae and neutron star mergers, potentially affecting energy transport and other processes. The simulation of fast flavor conversions under realistic astrophysical conditions requires substantial computational resources and involves significant analytical challenges. While machine learning methods like Multilayer Perceptrons have been used to accurately predict the asymptotic outcomes of FFCs, their ‘black-box’ nature limits the extraction of direct physical insight. To address this limitation, we employ two distinct interpretable machine learning frameworks—Kolmogorov-Arnold Networks (KANs) and Sparse Identification of Nonlinear Dynamics (SINDy)—to derive physically insightful models from a FFC simulation dataset. Our analysis reveals a fundamental trade-off between predictive accuracy and model simplicity. The KANs demonstrates high fidelity in reconstructing post-conversion neutrino energy spectra, achieving accuracies of up to 90%. In contrast, SINDy uncovers a remarkably concise, low-rank set of governing equations, offering maximum interpretability but with lower predictive accuracy. Critically, by analyzing the interpretable model, we identify that the system’s evolution is most sensitive to the initial number density of heavy-lepton neutrinos when FCCs are triggered, compared to other physical quantities. Ultimately, this work provides a methodological framework for interpretable machine learning that supports genuine data-driven physical discovery in astronomy and astrophysics, going beyond prediction alone.
@@ I Introduction
A significant recent advancement in this area is the discovery of fast flavor conversions (FFCs), which occur over remarkably short spatial scales in the vacuum [18, 19, 20, 21, 22, 23, 24]. A necessary and sufficient condition for the occurrence of FFC is that the angular distribution of the neutrino lepton number, defined as:
$G(\mathbf{v})=\frac{\sqrt{2}G_{\text{F}}}{(2\pi)^{3}}\int E_{\nu}^{2}dE_{\nu}\Big[(f_{\nu_{e}}(\mathbf{p})-f_{\nu_{x}}(\mathbf{p}))$ (1)
$-(f_{\bar{\nu}_{e}}(\mathbf{p})-f_{\bar{\nu}_{x}}(\mathbf{p}))\Big],$
[…]
The extremely short characteristic scales of FFC, in stark contrast to the much coarser resolution of typical hydrodynamic simulations of CCSNe and NSMs, present a substantial challenge for seamless integration. The vast disparity in temporal and spatial scales across the various physical processes further renders direct, first-principles simulations computationally prohibitive. Compounding these difficulties, certain simulation details also require careful attention—for instance, while the equipartition condition appears reasonable for out-going neutrinos, significant deviations arise for incoming neutrinos, affecting the matter profiles [26]. A promising strategy is to conduct small-scale FFCs simulations and apply the insights gained to inform larger hydrodynamic models [27, 28, 29, 30, 31, 32, 33]. To address this, extensive research has focused on local dynamical simulations with periodic boundary conditions [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], where quasistationary flavor states have been observed in the neutrino gas and are well described by analytical models [45]. Recent studies suggest that when the initial symmetry of the neutrino field is broken, the nature of flavor conversion changes, potentially offering deeper insights into FFCs [46, 23]. On the theoretical side, a quantum framework describing neutrino interactions with flavor wave quanta has recently been proposed [47]. On the computational side, an approximate method to predict the asymptotic state of FFCs without solving the full quantum kinetic equations in the three-flavor framework has also been developed [48]. These advances offer potential solutions to the computational challenges of fast flavor evolution.
@@
$I_{n,i}=\frac{E_{\nu,i}^{2}\Delta E_{\nu,i}}{(2\pi)^{3}}\int_{-1}^{1}d\mu\,\mu^{n}\int_{0}^{2\pi}d\phi_{\nu}\,f_{\nu}(p).$ (2)
Although flavor equilibration may occur immediately following rapid FFCs, this is not generally the asymptotic outcome, as feedback from subsequent collisional neutrino-matter interactions becomes increasingly important—particularly when FFCs occur inside the neutrinosphere [49]. Effective analytical methods have been developed to enable efficient and robust modeling of collisional feedback within the neutrinosphere [50]. Nonetheless, leveraging artificial intelligence to uncover hidden physical patterns in data remains a highly promising direction. This is because traditional analytical approaches typically involve specific assumptions and may not fully capture all aspects of the system’s complexity, creating unique opportunities for data-driven methods. With the rapid development of various neutrino telescopes, employing AI techniques offers an effective way to handle the vast volumes of observational data expected in the future and to extract the underlying physical insights. Recently, advances in machine learning, particularly neural networks, have been increasingly applied to astrophysics, including cosmological simulations [51, 52, 53, 54] and astronomical data processing [55, 56, 57, 58, 59, 60, 61, 62, 63]. Recent studies have shown that artificial neural networks (NNs) can accurately predict the asymptotic outcomes of FFCs in both single- and multi-energy neutrino gases within the moments framework [64, 65, 66, 67].
While deep learning has achieved remarkable success in certain research areas, many networks are parameter-intensive and lack transparency, making their internal mechanisms difficult to interpret. This remains a key challenge in AI-driven scientific research [68, 69]. This issue is often framed as a direct trade-off between accuracy and interpretability [70]. This trade-off is crucial when dealing with data that may contain complex, multivariate, and coupled features such as FFCs. Addressing this process requires moving beyond traditional black-box models, because a deeper understanding of the phenomena demands interpretability [71]. Interpretable networks are not a compromise for lower accuracy; they are the preferred outcome of a more rigorous model-selection process [72]. Interpretability enhances efficiency and, more importantly, enables the extraction of physical features by making a model’s decision basis traceable. By linking learned patterns to physical principles, interpretable models transform predictive tools into engines of discovery, potentially yielding new hypotheses and insights in astronomy and physics [73, 74].
Recently, Kolmogorov-Arnold Networks (KANs) have emerged as a promising solution. Inspired by the Kolmogorov-Arnold representation theorem, KANs introduce learnable activation functions that significantly improve model expressiveness while maintaining interpretability [75]. Compared to conventional neural networks that focus on extracting features from data, KANs are more inclined to discover underlying physical laws. In parallel, the Sparse Identification of Nonlinear Dynamics (SINDy) framework has demonstrated effectiveness in recovering compact, symbolic representations of governing equations from data [76]. By promoting sparsity in the candidate function space, SINDy identifies the minimal set of terms required to describe system dynamics, offering a powerful tool for discovering interpretable physical laws from simulation outputs or observations.
In section II, we briefly introduce the structure of the model input. section III details the construction of KANs. In section IV, we describe the working principles of SINDy and how it is applied to the FFCs dataset. The reconstruction performance of KANs and SINDy for post-FFC supernova neutrino spectra is demonstrated in section V. Subsequently, section VI presents the extraction of physical insights from both the trained KANs and SINDy models. In section VII, we discuss the trade-off between predictive accuracy and interpretability in KAN and SINDy. Finally, section VIII provides a summary of this work.
II Model Input
We first briefly introduce the input dataset used in this work, which is identical to that of [65] and follows the fast–flavor–conversion (FFC) framework of [36]. The data are produced by a 1D-box simulation with a periodic boundary in one spatial direction and translational symmetry in the other two, evolved for several thousand interaction timescales. It spans a grid of initial $\bar{\nu}_{e}/\nu_{e}$ number–density ratios and small seed perturbations, in the two–flavor, forward–scattering limit (neglecting momentum–changing collisions). Although simulations can realize FFCs under simplified conditions, they remain computationally expensive and difficult to embed directly into large-scale CCSN models. A surrogate model offers a physics-aware but lightweight framework that can be integrated into future simulations, while also enabling the extraction of physical insight from observational data or more complex numerical results. Rather than replacing the simulations, the surrogate serves as a bridge between detailed calculations and realistic astrophysical applications. For training and testing our models, the dataset we used combines maximum entropy and Gaussian initial neutrino angular distributions. The final outcome of FFCs is determined by the two-flavor survival probability:
$P_{\text{sur}}(\mu)=\begin{cases}\frac{1}{2},&\text{for }\mu^{<}\\ \mathcal{S}_{2}(\mu),&\text{for }\mu^{>}\end{cases},$ (3)
@@ II Model Input
where $\Theta$ denotes the Heaviside theta function, $G(\mu)=\int_{0}^{2\pi}\mathrm{~d}\phi_{\nu}G(\mathbf{v})$ [45, 64].For convenience, we refer to the interval of $\mu$ that makes the above integral small (large) as the ”small” (”large”) side, which corresponds to $\mu^{<}$ ($\mu^{>}$) in Equation 3
@@
Our 11-dimensional input vector consists of two sets of features. The first nine quantities primarily describe the moments of the neutrino distribution, as defined in Table 1. To keep the model general, in contrast to [65] that treats energy bins independently, we include two energy-dependent components, $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$, defined for each bin $i$ as follows [65]:
$E_{\nu,i}\;=\;70\;-\;60\,\sqrt{\frac{F_{\nu,i}}{F_{\nu}^{\rm tot}}}\,,\quad E_{\nu,i}\in[0,\,60],$ (5)
with an identical expression for antineutrinos. In practice, each $E_{\nu,i}$ is computed from the ratio $F_{\nu,i}/F_{\nu}^{\rm tot}$ and then clipped to the interval [0, 60][0,\,60] to ensure numerical stability. A mask is applied to retain only those bins for which both $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$ are nonnegative; this masking is applied consistently to both inputs and targets so that the original training‐set composition is preserved. Finally, $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$ are concatenated (in that order) to the nine moment‐integrated inputs, yielding an 11‐dimensional vector for each sample. By appending $E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$ as continuous coordinates of energy, a single unified network can condition its prediction on energy and thereby act as a continuous surrogate across energy, rather than as a per–bin regressor. Although in the FFCs process the learned dependence on $E$ is expected to be weak, including these tags is structurally important to maintain generality and portability when subleading, energy–dependent effects become non–negligible (see section VI for the energy–dependence discussion).
The model parameters are as follows: $\alpha\in(0,2.5)$, $F_{\nu_{x},(i)}\in(0,1)$, $\bar{F}_{\nu_{e},(i)}\in(0.4F_{\nu_{x},(i)},F_{\nu_{x},(i)})$, and $F_{\nu_{e},(i)}\in(0.4\bar{F}_{\nu_{e},(i)},\bar{F}_{\nu_{e},(i)})$, which follows the expected hierarchy: $F_{\nu_{e}}\lesssim\bar{F}_{\nu_{e}}\lesssim F_{\nu_{x}}$. The neutrino number densities $n_{\nu,i}$ are sampled from a half-normal distribution with a mean of zero and a standard deviation of $0.1n_{\nu}$, where $n_{\nu}$ is the energy-integrated neutrino number density. With $F_{\nu}=(I_{1}/I_{0})_{\nu}$ and $\alpha=n_{\bar{\nu}_{e}}/n_{\nu_{e}}$, the quantities with and without the subscript $i$ refer to the energy-integrated spectrum and specific energy bin, respectively. The neutrino number densities in a given energy bin must be smaller than the corresponding energy-integrated values. This is consistent with [65]. Moreover, the dataset used for both modeling approaches is consistent with [65], including the sample size, the train/validation split for the KAN network in section III, and other detailed settings. In particular, the SINDy method in section IV does not require a train/validation split and is fitted directly on the full dataset. Because the core of SINDy is to use sparse regression to select as few non-zero coefficients as possible from the candidate library, resulting in the ”simplest” control equations; this step itself is a structural selection and does not depend on training/validation splitting.
@@ III Kolmogorov-Arnold Networks
$\color[rgb]{0,0,0}f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$ (6)
@@
$\color[rgb]{0,0,0}\text{KANs}(x)=(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{0})(x).$ (8)
@@
We also define two terms used to assess model performance. Both metrics are forms of the Mean Absolute Error (MAE) computed on the validation set. A small constant $\epsilon=10^{-8}$ is added to the denominator to ensure numerical stability.
The overall relative error, denoted $\mathcal{L}_{\text{rel}}$, measures the average relative error across all eight output neurons. For a validation set with $N$ samples, where $\hat{y}_{j,i}$ is the $i$-th predicted output for sample $j$ and $y_{j,i}$ is the corresponding true value,
The $\bm{\nu_{e}+\bar{\nu}_{e}}$ error, denoted $\mathcal{L}_{\nu}$, evaluates the model’s ability to predict the total contribution of electron neutrinos and antineutrinos. For each sample $j$, let $S_{\text{pred},j}$ and $S_{\text{true},j}$ be the predicted and true totals, respectively. The loss is
Figure 1 then reports these metrics over the course of training: the blue curve corresponds to $\mathcal{L}_{\text{rel}}$ and the red curve to $\mathcal{L}_{\nu}$.
@@
Figure 2: Visualization of the trained KAN architecture used to model the fast flavor conversion process. The eleven input neurons at the bottom correspond, from left to right, to the physical features of the pre-conversion neutrino gas as detailed in Equation 4. The eight output neurons at the top yield the predicted post-conversion radial moments for all relevant flavors ($I_{0}^{\nu_{e}},I_{1}^{\nu_{e}},I_{0}^{\bar{\nu}_{e}},I_{1}^{\bar{\nu}_{e}},I_{0}^{\nu_{x}},I_{1}^{\nu_{x}},I_{0}^{\bar{\nu}_{x}},I_{1}^{\bar{\nu}_{x}}$, from left to right). The curves on the edges represent the learned one-dimensional activation functions, which are parameterized by B-splines. The activation shapes shown are taken from a representative training run, illustrating the complex functional relationships captured by the model. Edge darkness denotes the contribution of the connected input (darker means higher contribution); nearly invisible edges indicate negligible contribution.
In the KAN framework, each learnable activation is parameterized by B-splines, which can subsequently be approximated by elementary analytic functions through the built-in symbolic regression module. This procedure enhances interpretability by providing symbolic results of the learned nonlinearities without requiring any modification of the original KAN implementation. In our study, we used the default symbolic regression functionality included in the KAN library [75].
However, the symbolic regression process exhibits a notable sensitivity to the random seed used in training. To obtain a robust and canonical symbolic model, we adopted an ensemble averaging approach. Specifically, the network was trained multiple times with different random seeds, and the B-spline coefficients for each learnable activation function were averaged across all runs. By performing symbolic regression on this stable set of averaged parameters, we mitigate the effects of stochastic initialization and derive a more representative model of the underlying dynamics.